OpenStudy (anonymous):

The curve y=ax^2 + b passes A (3,39) and B (-2,c). The tangent to the curve at A is parallel to the line y=30x-7. find a,b and c and the tangent to the curve at B

OpenStudy (anonymous):

Oh this is fun!

OpenStudy (ash2326):

@Lachlan1996 Do you know calculus?

OpenStudy (anonymous):

If you take the derivative of ax^2+b, what do you get?

OpenStudy (anonymous):

yes, did it months and months ago, been on holidays now

OpenStudy (anonymous):

uhhh yes but im not sure if b contains an x so you may not be able to remove b when you derive to get 2ax

OpenStudy (anonymous):

I think that you are overthinking it... treat b like a constant.

OpenStudy (anonymous):

so... y'=2ax... the derivative is a "slope function", so steal the slope from the equation for the tangent line that they give you.

OpenStudy (ash2326):

Yes, a and b are just constants

OpenStudy (anonymous):

so 30 = 2a(3)

OpenStudy (anonymous):

awesome! solve for a

OpenStudy (anonymous):

a = 5

OpenStudy (anonymous):

Now use that in the original to solve for b given the point (3,39)

OpenStudy (anonymous):

ahh yes so then 39 = a(3)^2 + b

OpenStudy (anonymous):

thus 39 = 9a + b

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

39 = 9(5) + b 39 = 45 + b b = -6

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

now use everything that you know with that last point and the original equation.

OpenStudy (anonymous):

okie dokie, therefore to find c y=5x^2 -6 = 5(-2)^2 -6

OpenStudy (anonymous):

therefore c =14

OpenStudy (anonymous):

fantastic! It was like a puzzle... how fun!

OpenStudy (anonymous):

and gradient of B would be y' = 10x therefore y'= 10(-2) = -20 thus y=mx+c -2=-20(-2)+c

OpenStudy (anonymous):

no wait sorry 14 = -20(-2) + c

OpenStudy (anonymous):

therefore c=-26 thus the equation for the tangent at B would be y=-20x-26

OpenStudy (anonymous):

Thankyou guys you are awesome!

OpenStudy (anonymous):

um... a=5, b=-6, and c=14... done!

OpenStudy (anonymous):

oh yeah, tangent to the curve at B... y'=2ax=2(5)(-2)=-20

OpenStudy (anonymous):

Thankyou for the help mate :) have a great day

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

alrighty cheers for that mate, ill close the question now, and i wish you well